Do people take indoor swimming lessons in the middle of the hot summer? They sure do at the Webster Aquatic Center. During the month of July 2009 alone, 179 people participated in various forms of lessons.
Swim Categories Daytime Evenings
Preschool 27 33
Levels 76 38
Adult and diving 4 6
Total 107 77
Suppose one swimmer was selected at random from the July participants.
(d) Find P(preschool).
(e) Find P(daytime).
(f) Find P(not levels).

(g) Find P(preschool or evening).

(h) Find P(preschool and daytime).

(i) Find P(daytime | levels).
(j) Find P(adult and diving | evening).

Answer by stanbon(75887)

Do people take indoor swimming lessons in the middle of the hot summer? They sure do at the Webster Aquatic Center. During the month of July 2009 alone, 179 people participated in various forms of lessons.
Swim Categories Daytime Evenings Totals
Preschool 27 33 60
Levels 76 38 114
Adult and diving 4 6 10
Total 107 77 184
-----------------------------------------
Suppose one swimmer was selected at random from the July participants.
(d) Find P(preschool) = 60/184
(e) Find P(daytime) = 107/184
(f) Find P(not levels) = (60+10)/184
(g) Find P(preschool or evening) = (60+77-33)/184
(h) Find P(preschool and daytime) = 27/184
(i) Find P(daytime | levels) = P(day AND levels)/P(levels) = 76/114
(j) Find P(adult and diving | evening) = 6/77
============
Cheers,
Stan H.
------------